OpenStudy (anonymous):

point P is at (2,4) and point Q is at (3,4).find PQ

OpenStudy (saifoo.khan):

Use the Distance Formual.

OpenStudy (saifoo.khan):

Formula*

OpenStudy (anonymous):

Is PQ referring to the distance or the line segment connecting P and Q?

OpenStudy (anonymous):

i think distance

OpenStudy (anonymous):

Its obviously 1, just draw those two points in a plane!

OpenStudy (anonymous):

No need for distance formula either.

jimthompson5910 (jim_thompson5910):

You can use the distance formula d = sqrt((x2-x1)^2+(y2-y1)^2) In this case, x1=2, y1=4, x2=3, and y2=4. Remember that the two generic points are (x1, y1) and (x2, y2) So (x1, y1) = (2, 4), which means x1=2 and y1=4 and (x2, y2) = (3, 4), which means x2=3 and y2=4 ------------------- d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((3-2)^2+(4-4)^2) d = sqrt((1)^2+(0)^2) d = sqrt(1+0) d = sqrt(1) d = 1 So exact distance between the two points is 1 unit.

jimthompson5910 (jim_thompson5910):

Or you can take the shortcut by noticing that they two points have the same y coordinate and just subtracting the x coordinates to find your answer

OpenStudy (anonymous):

thnx